the p term of an ap is q and q term is p .find its p+q term 

Grade:11

1 Answers

Arun
5 years ago
 

The pth term of an AP is = a+(p-1)d

 

where a is the first term of the AP and  d is the common difference

 

as pth term is equal to q therefore q=a+(p-1)d .....1

 

similarly qth term is equal to p

 

p=a+(q-1)d .......2

 

solve the two equations to get the value of a and d. It will come

 

a=p+q-1

 

d=-1

 

now (p+q)th term is =a+(p+q-1)d

 

put the values of a and d

 

=(p+q-1)+(p+q-1)(-1)

 

=(p+q-1)-(p+q-1)

 

=0